OpenStudy (kaydra14):

The graph of the function f(x) = 5^x - 12 contains the point (2 , y). What is the value of y ?

OpenStudy (jack1):

f(x) = 5^x - 12 (x,y) so at the point x = 3, y = ? (3,y) f(x) = 5^x - 12 f(3) = 5^3 - 12 f(3) = 125 - 12 f(3) = 113 so y = 113 when x = 3 point is (3,113) now u try, if x = 2 (2,y) ==> ????

OpenStudy (jack1):

@kaydra14 ...u there?

OpenStudy (jack1):

are u ok with the above example... do u need anything explained more?

OpenStudy (jack1):

kaydra how are you going with this?

OpenStudy (kaydra14):

no i got , i am working it out now

OpenStudy (jack1):

sweet ;)

OpenStudy (kaydra14):

thank you

OpenStudy (jack1):

np c: what did u get?

OpenStudy (kaydra14):

is the right answer ( 2,13 )

OpenStudy (jack1):

yep!